OpenStudy (anonymous):

2. What is the equation, in standard form, of a parabola that models the values in the table?

OpenStudy (anonymous):

x -2 0 4 f(x) 0 -6 78

OpenStudy (anonymous):

a. y=6x^2+5x-4 b. y=-4x^2-5x+6 c. y=5x^2+4x-6 d. y=4x^2+5x-6